OpenStudy (anonymous):

how do you get the derivative of ln(tanx)? because i got 1/tanx*sec^2x but it's wrong?

OpenStudy (campbell_st):

well isn't tan = sin/cos and I think you need to look at the log laws so you are looking at \[ \ln \frac{\sin(x)}{\cos(x)} = \ln(\sin(x)) - \ln(\cos(x))\] now you can take the derivative

OpenStudy (cwrw238):

or you use the chain rule derivative = (1/ tanx ) * sec^2x

OpenStudy (cwrw238):

= 1 cos x ---- * ---- = 1 / cosx sinx cos^2 sinx is one way of writing it

OpenStudy (xapproachesinfinity):

the derivative of that just take the inside derivative which is 1+tan^2x then the derivative of the outer function which is 1/tanx so \(\Large {\frac{1+tan^2x}{tanx}}=\normalsize cotx+tanx\)

OpenStudy (xapproachesinfinity):

derivative of tanx is (secx)^2 which is also 1+(tanx)^2

OpenStudy (phi):

your answer is ok. There are a few ways to write it.

OpenStudy (anonymous):

thanks guys but the answer is 2/sin2x , because i have a maths book with all the answers written in the back of the book, just not the methods... i literally have no idea to go from the answer i got to that answer... help pls

OpenStudy (cwrw238):

yes - many ways

OpenStudy (phi):

sin(2x) = 2 sin(x) cos(x)

OpenStudy (xapproachesinfinity):

you can also use tanx=sinx/cosx like @campbell_st did Ate any rate just be aware that there are many ways to right trig functions using identities like @phi said

OpenStudy (phi):

and 2/(2 sin x cos x) = 1/(cos x sin x)

OpenStudy (cwrw238):

yes 2 / sin2x = 2 / 2sinx cosx = 1 / sinx cosx

OpenStudy (phi):

Here is a graph of (1/tan x) * sec^2(x) and 2/sin(2x)

OpenStudy (xapproachesinfinity):

@phi may i ask what graphing web you used?

OpenStudy (phi):

geogebra. It's free. google for it.

OpenStudy (xapproachesinfinity):

oh ok! thanks

OpenStudy (anonymous):

ok so now i wrote it as 1+tan^2x/tanx ... now what?

OpenStudy (xapproachesinfinity):

you can leave is sec^2x/tanx or put in any other different form you like there are all correct

OpenStudy (anonymous):

i just want to know how to get to 2/sin2x ,i'm really confused lol sorry i'm not that good at maths and we've just started differentiating 2 days ago

OpenStudy (freckles):

\[\frac{\sec^2(x)}{\tan(x)} =\frac{1}{\cos^2(x)} \cdot \frac{\cos(x)}{\sin(x)}\] But recall 2sin(x)cos(x)=sin(2x) so sin(x)cos(x)=sin(2x)/2

OpenStudy (anonymous):

@freckles where did you get 1/cos^\[1/\cos^2x . cosx/sinx ?\]

OpenStudy (freckles):

sec^2(x)=1/cos^2(x)

OpenStudy (freckles):

1/tan(x)=cos(x)/sin(x)

OpenStudy (anonymous):

@freckles how is sec^2x equal to 1/cos^2x? and how does tan become 1/tan(x) and how does 1/tanx become cosx/sinx?? sorry it doesn't make sense to me... and when u multiply cos^2x by sinx, where does the 2 go?

OpenStudy (freckles):

\[\frac{\sec^2(x) \cdot 1}{1 \cdot \tan(x)}=\frac{\sec^2(x)}{1} \cdot \frac{1}{\tan(x)}\]

OpenStudy (freckles):

\[\sec(x)=\frac{1}{\cos(x)} \\ \text{ squaring both sides gives } \\ (\sec(x))^2=(\frac{1}{\cos(x)})^2 \\ (\sec(x))^2=\frac{1^2}{(\cos(x))^2} \\ (\sec(x))^2=\frac{1}{(\cos(x))^2} \\ \sec^2(x)=\frac{1}{\cos^2(x)}\]

OpenStudy (freckles):

\[\frac{1}{\tan(x)}=\cot(x)=\frac{\cos(x)}{\sin(x)}\]

OpenStudy (freckles):

\[\frac{\sec^2(x)}{\tan(x)}=\sec^2(x)\cot(x)=\frac{1}{\cos^2(x)} \frac{\cos(x)}{\sin(x)} \\ =\frac{\cos(x)}{\cos(x)\cos(x)\sin(x)} \\ =\frac{ \cancel {\cos(x)}}{\cancel{\cos(x)} \cos(x)\sin(x)}\]

OpenStudy (freckles):

\[=\frac{1}{\frac{1}{2} 2 \sin(x)\cos(x)}\] You should be able to finish this...

OpenStudy (freckles):

given sin(2x)=2sin(x)cos(x)

OpenStudy (freckles):

I take it you didn't take trigonometry before calculus ?

OpenStudy (freckles):

This means you are going to have learn trig on your own (or with the help of OpenStudy) without the instructor going over the trig like stuff.

OpenStudy (freckles):

 \[\sin(\theta)=\frac{y}{r} \\ \cos(\theta)=\frac{x}{r} \\ \tan(\theta)=\frac{y}{x}=\frac{\frac{y}{r}}{\frac{x}{r}}=\frac{\sin(\theta)}{\cos(\theta)} \\ \csc(\theta)=\frac{r}{y}=\frac{1}{\frac{y}{r}}=\frac{1}{\sin(\theta)} \\ \sec(\theta)=\frac{r}{x}=\frac{1}{\frac{x}{r}}=\frac{1}{\cos(\theta)} \\ \cot(\theta)=\frac{x}{y}=\frac{\frac{x}{r}}{\frac{y}{r}}=\frac{\cos(\theta)}{\sin(\theta)}\]